OpenStudy (madhatter):

Find the slope of the line 5,1 4,5

OpenStudy (anonymous):

Alright, the first step is identifying y1, y2, x1, and x2. Do you know which is which? Also, do you know the equation for finding slope?

OpenStudy (madhatter):

yes x=5 y=1 then 4=z y=5 i dont not know the equation for finding slope

OpenStudy (anonymous):

the actual equation to start with is y=mx+b

OpenStudy (anonymous):

\[m=\frac{ y1-y2 }{ x1-x2 }\]

OpenStudy (madhatter):

ok good to know

OpenStudy (anonymous):

(5,1), (1,4) slope of the line is=(4-1)/(1-5) or slope=3/-4

OpenStudy (anonymous):

if you have \[(x _{1},y _{1}) and (x _{2},y _{2})\] and you are finding the slope, \[m=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\] where m is the slope....

OpenStudy (madhatter):

oh ok thank

OpenStudy (anonymous):

m is slope of the line, remember the slope equation, you'll use is a lot in the future. :) If you need the equation of the line, just plug your slope (m) and your y1 and x1 coords into the following equation and solve for y: y-y1=m(x-x1)

OpenStudy (madhatter):

ok thanks @melbel